OpenStudy (anonymous):

Solve the system of equations by using the substitution method. x^2+y^2=25 y-5= -x^2

OpenStudy (arbershabani97):

x^2+y^2=25 x^2+y=5 then y^2-y+20=0 (are you sure that you wrote it correctly) because you need to use quadratic formula there)

OpenStudy (anonymous):

Okay I got that far and then I was lost. Probably because nowhere in the lesson does it say to use quadratic formula..

OpenStudy (arbershabani97):

well probably it should've been x^2+y^2=-25 and it would make sense

OpenStudy (anonymous):

I know but its a positive. I copied the problem directly from the page.

OpenStudy (arbershabani97):

either technical problem, or given in purpose, but because you didn't learn the quadratic formula, then you'll probably learn it later, so you won't need it now, so don't worry about it

OpenStudy (radar):